# Cramming 1568 Tokens into a Single Vector and Back Again: Exploring the Limits of Embedding Space Capacity

### v2

@@ 2 Related Work
Another line of work targets compressing only the prompts to reduce inference costs. Gist tokens Mu et al. (2023) are prompt representations compressed by the LLM itself, finetuned with a special mask. Gisting allows to achieve prompt compression rate up to x26 with only minor loss in model performance. LLMLingua Jiang et al. (2023) decouples the compression operation from the LLM and introduces a coarse-to-fine prompt compression strategy with a budget controller and token-level iterative compression, achieving up to x20 compression with negligible performance loss. Jiang et al. (2024); Pan et al. (2024) extend this framework to long contexts, improving information retention through data distillation. Additionally, Morris et al. (2024) suggest that some information about prompts can be recovered from the model predictions themselves. In the current work we apply a per-sample optimization process instead to explore the fundamental limits of compression and establish upper bounds on compression rates that far exceed prior work.
@@ 4.5 Embedding Capacity Utilization
For models within the Llama family, we observe that the theoretical capacity is identical for both the 3B and 8B variants. However, in practice, the 8B model exhibits significantly higher capacity utilization. This might be fully attributed to better language understanding, gained by the larger model during pre-training, but capacity utilization for random text is higher for 8B model as well. This result suggests that the overall number of parameters plays an important role in determining effective capacity not only via LM capabilities but also due to better utilization of embedding space for episodic information storage.
@@ 4.6 Non-Transformer Models: Mamba
Notably, the findings from the previous sections are not limited to Transformer-based language models. State-space Mamba models can also encode an entire text into a single [mem] vector and perfectly reconstruct the text when its cross-entropy falls below the model’s capacity threshold, e.g., Mamba-1.4B can encode texts of $512$ tokens into a single [mem] vector, similar to other 1B models (LLama-3.2-1B, Sheared-LLaMa-1.3B). This demonstrates that obtained results are architecture-agnostic (more details on Mamba results in Appendices C, 6 and 3).
On capacity utilization (Fig. 5), the Mamba models exhibit a consistent pattern across both panels that depends on model size. In the top plot, the increase in correctly decoded tokens is nearly linear with respect to each model’s theoretical capacity. This implies that each additional parameter in the embedding space translates into greater decoding accuracy at a nearly constant rate. In the bottom plot, capacity utilization rises monotonically from the 130M to the 1.4B model. This is the opposite of the downward trend seen in Pythia and shows that larger Mambas progressively make better use of their learned input space.
@@ Appendix A Models and Training Details
Llama-3.2-3B meta-llama/Llama-3.2-3B 3.0 4096 128256

### v3

@@ 2 Related Work
Another line of work targets prompts compression to reduce inference costs. Gist tokens Mu et al. (2023) are prompt representations compressed by the LLM itself, finetuned with a special mask. Gisting allows to achieve prompt compression rate up to x26 with only minor loss in model performance. LLMLingua Jiang et al. (2023) decouples the compression operation from the LLM and introduces a coarse-to-fine prompt compression strategy with a budget controller and token-level iterative compression, achieving up to x20 compression with negligible performance loss. Jiang et al. (2024); Pan et al. (2024) extend this framework to long contexts, improving information retention through data distillation. Additionally, Morris et al. (2023, 2024) show that the original text can be reconstructed not only from its embeddings but even from the LM’s predictions. In the current work we apply a per-sample optimization process instead to explore the fundamental limits of compression and establish upper bounds on compression rates that far exceed prior work.
@@ 4.5 Embedding Capacity Utilization
For models within the Llama family, we observe that empirical capacity utilization rises steadily with model size (1B, 3B, 8B), most noticeable on natural texts from PG-19. This might be fully attributed to better language understanding, gained by the larger models during pre-training, but even on random texts we observe a modest upward trend. This result suggests that the overall number of parameters plays an important role in determining effective capacity not only via LM capabilities but also due to better utilization of embedding space for episodic information storage.
@@ 4.6 Non-Transformer Models: Mamba
Notably, the findings from the previous sections are not limited to Transformer-based language models. State-space Mamba models can also encode an entire text into a single [mem] vector and perfectly reconstruct the text when its cross-entropy falls below the model’s capacity threshold, e.g., Mamba-1.4B can encode texts of $512$ tokens into a single [mem] vector, similar to other 1B models (Llama-3.2-1B, Sheared-LLaMa-1.3B). This demonstrates that obtained results are architecture-agnostic.
On capacity utilization (Fig. 5), the Mamba models exhibit a consistent pattern across both panels that depends on model size. In the top plot, the increase in correctly decoded tokens is nearly linear with respect to each model’s theoretical capacity. This implies that each additional parameter in the embedding space translates into greater decoding accuracy at a nearly constant rate. In the bottom plot, capacity utilization rises monotonically from the 130M to the 1.4B model. This is the opposite of the downward trend seen in Pythia, but aligns with Llama, and shows that larger Mambas progressively make better use of their learned input space. We give more details on Mamba results in Appendices C, 6 and 3).
@@ Appendix A Models and Training Details
Llama-3.2-3B meta-llama/Llama-3.2-3B 3.0 3072 128256